OpenStudy (anonymous):

I need to reverse simplify this: 5y²-3y+6. Anyone have a good system for reverse FOILing, it always takes me forever ~_~.

OpenStudy (chaise):

Reverse FOILing is called factorising, and these types with co-efficients infront of the variable squared are a pain in the butt, and ontop of that, this graph doesn't cut the x axis and it really difficult to factorise

OpenStudy (anonymous):

Oh ok. Well i'm doing geometric sequences and trying to find the ratio so I can figure out the value of Y and according to the teachers notes this is the way to go about it and i'm comin up blank on the factorizing.

OpenStudy (kira_yamato):

It can't be factorized, because the discriminant < 0

OpenStudy (anonymous):

in my equation it is =0, I just didn't write it. So i did something wrong I guess. xD

OpenStudy (chaise):

Yes - that's right, the equation should be equal to zero. It must be equal to zero for it to be solved. I don't even know why your teacher would set anyone work to factorise this type of quadratic. The reason this can't be factorised is because b^2-4ac < 0. Therefore, in the quadratic formula, you cannot take the square root (as it is negative)

OpenStudy (anonymous):

Well this is what I started out with: Find Y so that 4y+1, y+4, 10-y are consecutive terms. then i went (y+4)/(4y+1) = (10-y)/(y+4) then cross multiplied and ended up with what you see in the above posts.

OpenStudy (kira_yamato):

The roots are imaginary... As you can see from the screenshot (drawn with Grapher)

OpenStudy (anonymous):

Consecutive terms in a geometric sequence**

OpenStudy (anonymous):

Is it solvable or is my teacher bonkers? lol

OpenStudy (chaise):

When you solve (y+4)/(4y+1) = (10-y)/(y+4) you do not get 5y²-3y+6. This could be your problem. Just a slight typo which has made an error in all our work. (y+4)/(4y+1) = (10-y)/(y+4) 5y^2-31y+6= 0 This would make more sense as this CAN BE factorised (as b^2-4ac > 0). Although I would use the quadratic formula as their is a coefficient infront of the y^2. I hope this has solved your dilemma. Are you aware of how to use the quadratic formula?

OpenStudy (anonymous):

omg i typed 3y ~_~ sigh. I did get 31y but i apparently am unable to type a 1 lol sorry.

OpenStudy (chaise):

It's fine - do you need any further help factorising 5y^2-31y+6= 0?

OpenStudy (anonymous):

And i'm not sure that I do. I'm taking a 30 level class after 10 years between when i took the pre-requisite and struggling to get my basics back.

OpenStudy (anonymous):

What is the quadratic formula?

OpenStudy (anonymous):

and yes i would seriously appreciate the help! haha

OpenStudy (kira_yamato):

For a quadratic equation \[ax^2 + bx + c\] \[x = \frac{-b ± \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

will i still get 2 values for x this way?

OpenStudy (anonymous):

you can try calculatersoup.com